A snowball is thrown with an initial x velocity of 7.5 m/s and an initial y velocity of 8.1 m/s .

KHW 4.A Lesson Check 4.48 5of 7> Constants Part A A snowball is thrown with an initial z velocity of 7.5 m/s and an initial y velocity of 8.1 m/s How much time is required for the snowball to reach its highest point? (Hint: The highest point of a projectile corresponds to the time when vy Express your answer using two significant figures and include the appropriate units. 0.) tValue Units Submit Previous Answers Request Answer incorrect; Try Again Provide Feedback Next >
Answer #1

The vertical motion of the ball is an uniformly accelerated motion, with constant acceleration a = -9.8m/s2(acceleration of gravity) is given by equation of motion v = u + at,

v (y direction) = 0m/s (since on reaching the maximum height final velocity will be zero

u( y direction) = 8.1m/s

a = -9.80665 m/s2

putting in given equation,

t = (v - u)/a

= 8.1/9.80665 (m/s) / (m/s2)

=0.825970 sec

= 0.83 sec (answer)